How many moles of barium metal react to produce 0.100mol of barium nitride? Hpw many moles of nitrogren gas react?

Ba(s) + N2(g) -> Ba3N2(s)
1
Expert's answer
2012-10-02T07:16:14-0400
x y & 0.100
3Ba(s) + 2N2(g) ------> Ba3N2(s)
3 & 2 1

x=3*0.100/1=0.300 mol

y=2*0.100/2-0.200 mol